There are 272 employees in an organization, 180 employees are graduates and 95 are females and 145 graduates are males. What is the probability that a randomly selected employee is;
(i) graduate and female (ii) male and not graduate.

Answer by Edwin McCravy(20063)

Make this chart, filling in what's given:

       | Graduates | Non-graduates || Totals |
-------|-----------|---------------||--------|
Males  |   145     |               ||        |
-------|-----------|---------------||--------|
Female |           |               ||   95   |
===================================||========|
Totals |   180     |               ||  272   |

We can fill in the total males by subtracting 272-95=177
We can fill in the total non-graduates subtracting 272-180=92
We can fill in the total female graduates subtracting 180-145=35

       | Graduates | Non-graduates || Totals |
-------|-----------|---------------||--------|
Males  |   145     |               ||  177   |
-------|-----------|---------------||--------|
Female |    35     |               ||   95   |
===================================||========|
Totals |   180     |      92       ||  272   |

Now we can fill in the total Male non-graduates by subtracting 177-145=32
We can fill in the total female non-graduates by subtracting 95-35=60


       | Graduates | Non-graduates || Totals |
-------|-----------|---------------||--------|
Males  |   145     |      32       ||  177   |
-------|-----------|---------------||--------|
Female |    35     |      60       ||   95   |
===================================||========|
Totals |   180     |      92       ||  272   |

Now that we have completed the chart we are ready to answer any probability
question they could ask us.
(i) graduate and female
That's 35 out of 272 or 35/272.
(ii) male and not graduate.
That's 32 out of 272 or 32/272 which reduces to 2/17.

Edwin